3 times the first of three consecutive numbers is 4 more than 2 times the second. Write an equation and solve the problem.
Answer by stanbon(75887)

3 times the first of three consecutive numbers is 4 more than 2 times the second. Write an equation and solve the problem.
---------------
1st: x
2nd: x+1
3rd: x+2
----------------
Equation:
3x = 2(x+1)+4
3x = 2x+2+4
x = 6 (1st)
x+1 = 7
x+2 = 8
==============
Cheers,
Stan H.